# Reference Points, Risk-Taking Behavior, and Competitive Outcomes in Sequential Settings

### v3

@@ 3.1 Lifting stage
Consider the attempt and lifting stages backward. We do not incorporate strategic decisions across stages to avoid complications making the situation intractable, but this would be an interesting research avenue. As noted earlier, the change in ranking takes values in $\{-1,0,1\}$, meaning that $\alpha$ does not affect this stage. Let $q(w_{H})$ denote the ex-ante probability that the higher-ranked rival successfully lifts the attempted weight $w_{H}$, which is unknown to lifter $i$ at the time of their own attempt. Similarly, let $q(w_{L})$ denote the ex-ante probability that the lower-ranked rival succeeds in lifting the attempted weight $w_{L}$, which is also unknown to lifter $i$ at the attempt stage.
Table 4: Utility Tuple $(v_{1}(x),(v_{2}(x),v_{3}(x)))$
Lower rival’s outcome turning around $i$ Higher rival’s attempt is turned around by $i$ No Yes No $(0,(0,0))$ $(0,(0,1))$ Yes $(-\lambda,(0,0))$ $(-\lambda,(0,1))$
Table 4 describes the utility tuple $(v_{1}(x),(v_{2}(x),v_{3}(x)))$ conditional on four scenarios determined by pressure from lower and higher rivals, where $v_{1}(x)$ is the utility when lifter $i$’s attempt weight $x$ fails, and $(v_{2}(x),v_{3}(x))$ is the utility when lifter $i$’s attempt is successful and the higher rival’s attempt either succeeds or fails. For simplicity, we assume $\alpha=1$, which does not affect our prediction. If there is no pressure from either lower or higher rivals (i.e., (No, No)), the current lifting attempt does not change lifter $i$’s rank whether successful or not, meaning $(v_{1}(x),(v_{2}(x),v_{3}(x)))=(0,(0,0))$. If there is pressure only from a lower rival (i.e., (Yes, No)), lifter $i$ experiences disutility $-\lambda(<-1)$ upon failure, but zero otherwise. If there is pressure only from a higher rival (i.e., (No, Yes)), lifter $i$ gains 1 if the higher rival fails, with the probability of $1-q(w_{H})$. If there is pressure from both lower and higher rivals, lifter $i$ experiences disutility $-\lambda(<-1)$ upon failure, gains 00 when both lifter $i$ and the higher rival succeed, and gains 1 if the higher rival fails, with the probability of $1-q(w_{H})$. Based on these scenarios, lifter $i$ has the strongest incentive to make an effort to avoid loss and gain when there is pressure from his lower and higher rivals. Based on these scenarios, lifter $i$ has the strongest incentive to exert effort to avoid loss and seek gains when under pressure from both lower- and higher-ranked rivals. If $\lambda>1$, the incentive to exert effort to avoid loss is further amplified.
3.2 Attempt stage
We now examine the attempt stage, focusing on how lifter $i$’s choice of attempt weight is influenced by their competitors. For simplicity, we analyze the effects of the lower and higher rivals separately. This distinction is justified by the empirical analysis, where both sources of pressure are incorporated simultaneously into the model. By conditioning on the influence of one competitor while examining the effect of the other, we can isolate their respective impacts on the lifter’s decision-making. [^10] While it is possible to construct a model that simultaneously accounts for both rivals, doing so significantly increases the complexity of the framework. To maintain analytical tractability and improve interpretability, we present the separate analyses in the main text. A more comprehensive model incorporating both rivals jointly is provided in Appendix A.1.
3.2.1 Response to a lower rival
Let $w_{L}$ denote the attempt weight of the lower rival and let $y_{L}$ and $y_{i}$ be the best outcome weights from earlier attempts within the competition for the lower rival and lifter $i$, respectively. The inequality conditions $y_{L}<y_{i}$, determined by rank order, and $y_{L}\leq w_{L}$ and $y_{i}\leq w_{i}$, meaning that the outcome weight must not exceed the corresponding attempt weight, must hold, given by the current competition rule. We consider three cases: [^11] If $w_{i}=w_{L}$, the lifter with the lighter bodyweight receives the higher ranking. Such tie-weight situations are categorized into the three cases described above.
Case 1: $w_{L}>w_{i}\geq y_{i}$. Even if lifter $i$ succeeds, his attempt weight $w_{i}$ remains below the lower rival’s $w_{L}$; thus, his rank depends solely on the lower rival’s success probability $q(w_{L})$. The expected utility is
$EU|_{w_{L}>w_{i}\geq y_{i}}=$ $\left[1-q(w_{L})\right]\cdot 0+q(w_{L})\cdot[-\lambda]$
$=$ $-\lambda q(w_{L})\leq 0.$
Case 2: $w_{i}>w_{L}>y_{i}$. In this case, if lifter $i$ succeeds his rank remains unchanged; however, if he fails and the lower rival succeeds, he incurs a loss. Thus,
$EU|_{w_{i}>w_{L}>y_{i}}=$ $q(w_{i})\cdot 0+\left[1-q(w_{i})\right]\left[1-q(w_{L})\right]\cdot 0+\left[1-q(w_{i})\right]q(w_{L})\cdot[-\lambda]$
$=$ $-\lambda\left[1-q(w_{i})\right]q(w_{L})\leq 0.$
Case 3: $w_{i}\geq y_{i}>w_{L}$. Here, even if the lower rival succeeds, there is no change in rank; hence,
$EU|_{w_{i}\geq y_{i}>w_{L}}=0.$
Since $w_{L}$ is exogenously determined from the viewpoint of lifter $i$, we obtain the ordering
$EU\mid_{w_{i}\geq y_{i}>w_{L}}\geq EU\mid_{w_{i}>w_{L}>y_{i}}\geq EU\mid_{w_{L}>w_{i}\geq y_{i}}.$
Moreover, note that
$\frac{\partial EU\mid_{w_{i}>w_{L}>y_{i}}}{\partial w_{i}}=\lambda q(w_{L})q^{\prime}(w_{i})\leq 0,$
since $q^{\prime}(w_{i})<0$. Therefore, lifter $i$ will declare an attempt weight that is slightly heavier than that of the lower rival. Note that the value of $\lambda$ does not directly affect the level of the chosen attempt weight. As long as $\lambda>0$, the prediction remains informative and aligned with observed behavior.
3.2.2 Response to a higher rival
Let $w_{H}$ denote the declared weight of the higher rival. Although this is not observable to lifter $i$’s attempt stage, we treat this as highly predictable information to lifter $i$ [^12] Using our data, we discuss the predictability of higher rival’s attempt weight given observable information at lifter $i$’s attempt stage in Section 7. Additionally, let $y_{H}$ be the higher rival’s best outcome weights up to the previous stage. The following cases arise:
Case 1: $w_{H}\geq y_{H}>w_{i}$. When lifter $i$’s attempt weight $w_{i}$ is below the higher rival’s outcome up to the previous stage, there is no change in rank, so
$EU|_{w_{H}\geq y_{H}>w_{i}}=0.$
Case 2: $w_{H}>w_{i}>y_{H}$. In this case, if lifer $i$ succeeds and the higher rival fails, he overtakes the higher rival; otherwise, his rank remains unchanged. Hence,
$EU|_{w_{H}>w_{i}>y_{H}}=$ $q(w_{i})\left[1-q(w_{H})\right]\cdot 1$
$=$ $q(w_{i})\left[1-q(w_{H})\right]\geq 0.$
Case 3: $w_{i}>w_{H}\geq y_{H}$. By declaring a weight higher than the higher rival’s, lifter $i$ secures the rank improvement if he succeeds. Thus,
$EU|_{w_{i}>w_{H}\geq y_{H}}=$ $q(w_{i})\cdot 1+\left[1-q(w_{i})\right]\cdot 0$
$=$ $q(w_{i})\geq 0.$
Let $w_{i}^{+}$ denote a choice such that $w_{i}>w_{H}$ and $w_{i}^{-}$ denote a choice with $w_{i}<w_{H}$. Comparing the expected utilities, we have:
$EU|_{w_{i}>w_{H}\geq y_{H}}>EU|_{w_{H}>w_{i}>y_{H}}\Leftrightarrow\frac{q(w_{i}^{+})}{q(w_{i}^{-})}>1-q(w_{H}),$
noting that $q(w_{i}^{+})<q(w_{i}^{-})$ since higher attempt weight implies a lower success probability and therefore $q(w_{i}^{+})/q(w_{i}^{-})\in(0,1)$.
Given $w_{H}$ is fixed, if the higher rival’s success probability $q(w_{H})$ is sufficiently high and/or if the success probability $q(w_{i})$ does not change significantly whether lifter $i$ chooses a weight that exceeds $w_{H}$ or one that remains below $w_{H}$, then we obtain
$EU|_{w_{i}>w_{H}\geq y_{H}}\geq EU|_{w_{H}>w_{i}>y_{H}}\geq EU|_{w_{H}\geq y_{H}>w_{i}}.$
Otherwise, if $q(w_{i})$ declines steeply when $w_{i}$ exceeds $w_{H}$, then the reverse ordering holds:
$EU|_{w_{H}>w_{i}>y_{H}}\geq EU|_{w_{i}>w_{H}\geq y_{H}}\geq EU|_{w_{H}\geq y_{H}>w_{i}}.$
Furthermore, note that
$\frac{\partial EU|_{w_{i}>w_{H}\geq y_{H}}}{\partial w_{i}}=q^{\prime}(w_{i})<0,\ \text{and}\ \frac{\partial EU|_{w_{H}>w_{i}>y_{H}}}{\partial w_{i}}=q^{\prime}(w_{i})\left[1-q(w_{H})\right]\leq 0.$
Therefore, lifter $i$ will declare a weight that is slightly heavier than either the higher rival’s (expected) declared weight or the higher rival’s best outcome up to the previous stage (i.e., $y_{H}$), depending on the higher rival’s success probability $q(w_{H})$ and on the sensitivity of $q(w_{i})$ to changes in $w_{i}$.
@@ 4.1 Choice of attempt weight
where $\tilde{W}_{it}^{k}$ is the difference between lifter $i$’s attempt weight compared to his current best in competition $t$ at a $k$-th attempt, $X_{it}$ is observed characteristics of lifter $i$ and competition $t$ including gender, body weight, number of competition experiences, a dummy variable for first participation, and fixed effects for competition $t$’s equipment category, age class, division, weight class, and federation. It also includes a dummy variable indicating whether the lifter shares the same declared attempt weight with another competitor, potentially causing ambiguity in the lifting order due to unresolved tie-breaking rules, and another dummy variable capturing mismatches between the actual lifting order and the interim ranking order, which may affect the set of observable information about rivals. $Z_{it}^{k}$ is a pressure variable during attempt defined later, $\varepsilon_{it}$ is assumed to be an i.i.d. error, and $\beta$ and $\gamma$ are vectors of coefficient parameters of $X_{it}$ and $Z_{it}^{k}$. Large $\tilde{W}_{it}^{k}$ implies his choosing a more challenging attempt weight. Our primary interest lies in $\gamma$, that is, the sensitivity to pressure.
The pressure variables, $Z_{it}^{k}$, which are exogenous due to the sequential game setting, represent the difference between the lifter’s current realized outcome and the attempted weight of the lower-ranked competitor at the next attempt, and the difference between the lifter’s current realized outcome and the expected attempt weight of the higher-ranked competitor predicted by the linear regression model. The actual higher-ranked attempt weight should not be used because it is not realized at lifter $i$’s attempt, although it must be predictable. We discuss the predictability issues and how to correct prediction errors, in particular, for constructing standard errors via bootstrap in Section 7. These variables capture the potential for the lifter’s rank to change by overtaking or being overtaken by the rivals.
@@ 4.2 Success probability
where $Y_{it}$ is whether the attempt is successful ($Y_{it}=1$) or not ($Y_{it}=0$), $\tilde{Z}_{it}^{k}$ is a pressure variable during lifting, $\eta_{it}$ is assumed to be an i.i.d. error, and $\delta$ represents the coefficient for the difference between lifter $i$’s attempt weight and his current best. Note that $\tilde{W}_{it}^{k}$, the choice of attempt weight, may be correlated with unobserved body conditions in $\eta$, potentially causing endogeneity issues. To address this, we employ a two-stage least squares (2SLS) regression, using Equation (1) as the first-stage regression, with the exogenous pressure variables during the attempt, $Z_{it}^{k}$, serving as instrumental variables (IV).
The pressure variable $\tilde{Z}_{it}^{k}$ during the lift is equal to one if the attempt involves overtaking a higher-ranked rival or being overtaken by a lower-ranked rival, and zero otherwise. These variables capture the potential for rank changes through overtaking or being overtaken.
@@ 5.1 Choice of attempt weight
Table 5: Regression of the Difference between the Second and Third Attempt Weight Compared to the Current Best on Pressure
@@
Pressure, lower rival, 2nd 0.106
Pressure, higher rival, 2nd 0.449
Pressure, lower rival, 3rd 0.057
Pressure, higher rival, 3rd 0.409
@@
Notes: We control gender, body weight, equipment category, age-class, division, weight-class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking. The standard errors in the brackets are clustered at federation categories. $W_{it}^{2}$ and $W_{it}^{3}$ are lifter $i$’s attempt weights in the second and third attempts in competition $t$. Pressure from lower rival at 2nd is (lower rival 2nd attempt) - (his 1st outcome). Pressure from lower rival at 3rd is (lower rival 3rd attempt) - (his 2nd outcome). Pressure from higher rival at 2nd is (higher rival 2nd attempt) - (his 1st outcome). Pressure from higher rival at 3rd is (higher rival 3rd attempt) - (his 2nd outcome).
Table 5 presents the regression results for the difference between the second and third attempt weights compared to the current best, incorporating various pressure variables. Male lifters exhibit mixed responses, with a statistically significant negative coefficient for the second attempt (–1.279) and a positive coefficient for the third attempt (3.583). This suggests that male lifters adopt a conservative approach in their second attempt but take greater risks in their third attempt, a pattern consistent with the literature on gender differences in risk-taking. Body weight and experience are positively associated with higher attempt weights in both attempts, indicating that heavier and more experienced lifters generally select heavier weights. The positive coefficients for first participation (32.122 and 33.623) suggest that first-time competitors attempt significantly higher weights relative to their current best.
@@ IV X X
Notes: We control gender, body weight, equipment category, age-class, division, weight-class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking. The standard errors in the brackets are clustered at federation categories.
@@
Overall, these findings highlight the role of pressure across attempts, with loss aversion playing a key role and overtaking higher-ranked rivals remaining particularly challenging. The asymmetry—where upward pressure consistently lowers success probabilities, while downward pressure has neutral or slightly positive effects—is consistent with the loss-averse utility structure assumed in the theoretical framework, whereby individuals exert more effort under pressure from lower-ranked rivals to avoid losses. [^13] This pattern is also observed in other competitive settings, such as marathons (Allen et al. 2017).

### v4

@@ 3.2 Lifting stage
We now turn to the lifting stage. In this stage, lifter $i$ chooses execution effort $e_{i}\geq 0$ to maximize expected utility, taking the declared attempt weight $w_{i}$ as given. Let $p_{i}(e_{i},w_{i})\in[0,1]$ denote lifter $i$’s probability of successfully lifting weight $w_{i}$. We assume $p_{i,e}(e_{i},w_{i})>0$ and $p_{i,w}(e_{i},w_{i})<0$, i.e., higher effort increases the success probability while heavier weights reduce it. Effort incurs a cost $C(e_{i})$ with $C^{\prime}(e_{i})>0$ and $C^{\prime\prime}(e_{i})>0$.
As in the attempt stage, interim rank serves as the reference point. The only payoff-relevant outcomes are whether xr=1x-r=1, xr=1x-r=-1, or xr=0x-r=0. The utility consequences of these outcomes are therefore given by the value function $v(\cdot)$ defined above.
We consider two benchmark situations. Under pressure from below (the risk of being overtaken), success prevents a rank loss while failure leads to a rank loss; expected utility is
$EU(e_{i};w_{i})=\bigl(1-p_{i}(e_{i},w_{i})\bigr)\cdot(-\lambda)-C(e_{i}).$
Under pressure from above, a rank gain occurs if $w_{i}>w_{H}$ and lifter $i$ succeeds, or if $y_{H}<w_{i}<w_{H}$ and lifter $i$ succeeds while the higher rival fails; otherwise the rank remains unchanged. Expected utility is therefore
$EU(e_{i};w_{i})=p_{i}(e_{i},w_{i})\bigl(\mathbf{1}(w_{i}>w_{H})+\mathbf{1}(y_{H}<w_{i}<w_{H})(1-q_{H}(w_{H}))\bigr)\cdot 1-C(e_{i}).$
In either case, the optimal effort $e_{i}^{*}$ satisfies the first-order condition
$p_{i,e}(e_{i}^{*},w_{i})\cdot V=C^{\prime}(e_{i}^{*}),$
where $V=\lambda>1$ under pressure from below and $V=\mathbf{1}(w_{i}>w_{H})+\mathbf{1}(y_{H}<w_{i}<w_{H})(1-q_{H}(w_{H}))\leq 1$ under pressure from above.
We further assume that the marginal effect of effort on success probability is weakly decreasing, i.e., $p_{i,ee}(e_{i},w_{i})\leq 0$, which captures diminishing returns to effort. Under this standard regularity condition and assuming an interior solution, the implicit function theorem implies that
$\frac{\partial e_{i}^{*}}{\partial V}=\frac{p_{i,e}(e_{i}^{*},w_{i})}{C^{\prime\prime}(e_{i}^{*})-V\,p_{i,ee}(e_{i}^{*},w_{i})}>0.$
Therefore, the possibility of a rank change increases optimal effort relative to a situation in which rank remains unchanged regardless of the outcome (i.e., $V=0$). Moreover, under loss-averse preferences ($\lambda>1$), effort responds more strongly to pressure from below than to pressure from above.
Hypothesis 3.
At the lifting stage, competitive pressure arising from potential rank changes increases execution effort and, consequently, the probability of success relative to situations in which rank is unaffected by the outcome. Furthermore, under loss-averse preferences, pressure from the lower rival (the risk of being overtaken) leads to a larger increase in success probability than pressure from the higher rival.
@@ 4.1 Choice of attempt weight
where $\tilde{W}_{it}^{k}$ is the difference between lifter $i$’s attempt weight and personal best as of attempt $k$ in competition $t$, and $X_{it}$ is a vector of observed characteristics of lifter $i$ and competition $t$, including gender, body weight, number of competition experiences, a dummy variable for first participation, and fixed effects for competition $t$’s equipment category, age class, division, weight class, and federation. It also includes a dummy variable indicating whether the lifter shares the same declared attempt weight with another competitor, potentially causing ambiguity in the lifting order due to unresolved tie-breaking rules, and another dummy variable capturing mismatches between the actual lifting order and the interim ranking order, which may affect the set of observable information about rivals. $Z_{it}^{k}$ is a pressure variable during the attempt stage, defined later; $\varepsilon_{it}$ is assumed to be an i.i.d. error; and $\beta$ and $\gamma$ are vectors of coefficients on $X_{it}$ and $Z_{it}^{k}$. A larger $\tilde{W}_{it}^{k}$ implies that the lifter chooses a more challenging attempt weight. Our primary interest lies in $\gamma$, that is, the sensitivity to pressure.
The pressure variables, $Z_{it}^{k}$, are exogenous due to the sequential game setting and are defined as rival-to-self gaps relative to the focal lifter’s current best outcome at the decision point. Specifically, pressure from the lower-ranked rival is measured as the lower rival’s declared attempt weight minus the focal lifter’s current best successful outcome, and pressure from the higher-ranked rival is measured as the expected higher rival attempt weight minus the focal lifter’s current best successful outcome. The actual higher-ranked attempt weight should not be used because it is not realized at lifter $i$’s attempt, although it must be predictable. We discuss the predictability issues and how to correct prediction errors, in particular, for constructing standard errors via the bootstrap in Section 7. These variables capture the potential for the lifter’s rank to change by overtaking or being overtaken by the rivals.
@@ 4.2 Success probability
where $Y_{it}$ is whether the attempt is successful ($Y_{it}=1$) or not ($Y_{it}=0$), $\tilde{Z}_{it}^{k}$ is a pressure variable during lifting, $\eta_{it}$ is assumed to be an i.i.d. error, and $\delta$ represents the coefficient for the difference between lifter $i$’s attempt weight and his personal best. Note that $\tilde{W}_{it}^{k}$, the choice of attempt weight, may be correlated with unobserved body conditions in $\eta$, potentially causing endogeneity issues. To address this, we employ a two-stage least squares (2SLS) regression, using Equation (1) as the first-stage regression, with the exogenous pressure variables during the attempt, $Z_{it}^{k}$, serving as instrumental variables (IV). Our IV strategy requires the exclusion restriction that, conditional on $X_{it}$ and $\tilde{Z}_{it}^{k}$, attempt-stage pressure $Z_{it}^{k}$ affects success probability $Y_{it}$ only through the chosen attempt weight $\tilde{W}_{it}^{k}$. This restriction is justified by two features of our design: (1) we control for lifting-stage pressure through $\tilde{Z}_{it}^{k}$, which captures realized rank-change pressure at execution, and (2) we intentionally model attempt and lifting stages separately, so any remaining direct channel from $Z_{it}^{k}$ to $Y_{it}$ is assumed to be limited after conditioning on $X_{it}$ and $\tilde{Z}_{it}^{k}$.
The pressure variable $\tilde{Z}_{it}^{k}$ during the lift consists of two indicator variables. The first, $\mathbf{1}(\text{Turned around})_{it}^{k}$, equals one if a lower-ranked rival’s $k$-th successful attempt exceeds lifter $i$’s current best outcome, and zero otherwise. When this indicator equals one, the lower rival has already recorded a lift above lifter $i$’s current best at that stage, so lifter $i$ would be overtaken in the interim ranking if he fails to improve upon his current best in the ongoing attempt. The second, $\mathbf{1}(\text{Turning around})_{it}^{k}$, equals one if lifter $i$’s $k$-th attempt weight exceeds the higher-ranked rival’s current best outcome, and zero otherwise. When this indicator equals one, a successful lift would place lifter $i$’s outcome above the higher rival’s current best, potentially reversing their ranking depending on the outcome of the higher rival’s subsequent attempt. These variables capture the potential for rank changes through being overtaken or overtaking, respectively.
@@ 5.1 Choice of attempt weight
Table 4: Regression of the Difference between the Second and Third Attempt Weight Compared to the Personal Best on Pressure
@@
Pressure, lower rival, 2nd 0.106***
Pressure, higher rival, 2nd 0.449***
Pressure, lower rival, 3rd 0.057***
Pressure, higher rival, 3rd 0.409***
@@
Notes: We control for gender, body weight, equipment category, age class, division, weight class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking. Standard errors in brackets are clustered at the federation level. $W_{it}^{2}$ and $W_{it}^{3}$ are lifter $i$’s attempt weights in the second and third attempts in competition $t$. Pressure from the lower rival at the second attempt is (lower rival second-attempt weight) - (focal lifter’s current best outcome after the first attempt). Pressure from the lower rival at the third attempt is (lower rival third-attempt weight) - (focal lifter’s current best outcome after the second attempt). Pressure from the higher rival at the second attempt is (predicted higher rival second-attempt weight) - (focal lifter’s current best outcome after the first attempt). Pressure from the higher rival at the third attempt is (predicted higher rival third-attempt weight) - (focal lifter’s current best outcome after the second attempt). $^{*}p<0.1$; $^{**}p<0.05$; $^{***}p<0.01$.
Table 4 presents the regression results for the difference between the second and third attempt weights compared to the personal best, incorporating various pressure variables. Male lifters exhibit mixed responses, with a statistically significant negative coefficient for the second attempt (–1.279) and a positive coefficient for the third attempt (3.583). This suggests that male lifters adopt a conservative approach in their second attempt but take greater risks in their third attempt, a pattern consistent with the literature on gender differences in risk-taking. Body weight and experience are positively associated with higher attempt weights in both attempts, indicating that heavier and more experienced lifters generally select heavier weights. The positive coefficients for first participation (32.122 and 33.623) suggest that first-time competitors attempt significantly higher weights relative to their current best.
@@ IV X X
Notes: We control for gender, body weight, equipment category, age class, division, weight class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking. Standard errors in brackets are clustered at the federation level. 1(Turned around) equals one if a lower-ranked rival’s successful attempt exceeds the lifter’s current best outcome, so that the lifter would be overtaken in the ranking if he fails to improve upon his current best in the ongoing attempt. 1(Turning around) equals one if the lifter’s attempt weight exceeds the higher-ranked rival’s current best outcome, so that a successful lift would place the lifter’s best outcome above that of the higher-ranked rival. $^{*}p<0.1$; $^{**}p<0.05$; $^{***}p<0.01$.
@@
Overall, these findings highlight the role of pressure across attempts, with patterns broadly consistent with loss-averse preferences and overtaking higher-ranked rivals remaining particularly challenging. The asymmetry—where upward pressure consistently lowers success probabilities, [^14] While the conceptual framework predicts that pressure from a higher-ranked rival increases effort, increased effort does not necessarily translate into higher success probabilities. In particular, choking under pressure may offset the effort response, resulting in lower observed success rates under upward pressure. while downward pressure has neutral or slightly positive effects—is consistent with the loss-averse utility structure assumed in the theoretical framework, whereby individuals exert more effort under pressure from lower-ranked rivals to avoid losses. [^15] This pattern is also observed in other competitive settings, such as marathons (Allen et al. 2017).